On December​ 1,2014​,the Itali Wholesale Co. is attempting to project cash receipts and disbursements through January​ 31,2015.

On this later​ date, a note will be payable in the amount of 107,000.

This amount was borrowed in September to carry the company through the seasonal peak in November and December.

Selected general ledger balances on December 1 are as follows:

Cash

$30,000

Inventory

111,800

Accounts payable

$139,000

Sales terms call for a 3% discount if payment is made within the first 10 days after the​ sale, with the balance due by the end of the month after sale. Experience has shown that 50​% of the billings will be collected within the discount period in the month of the​ purchase, 30​% by the end of the month after​ purchase, and 15​% in the following month. The remaining 5​% will be uncollectible. There are no cash sales.

The average selling price of the​ company's products is

$170

per unit. Actual and projected sales are as​ follows:

October actual

$287,000

November actual

629,000

December estimated

561,000

January estimated

612,000

February estimated

510,000

Total estimated for year ending June 30, 2015

$3,218,750

All purchases are payable within 15 days. Approximately 60​% of the purchases in a month are paid that month and the rest the following month. The average unit purchase cost is $130.Target ending inventories are 570 units plus 20​%of the next​ month's unit sales.

Total budgeted​ marketing, distribution, and​ customer-service costs for the year are $670,000. Of this​ amount, $155,000 are considered fixed​ (and include depreciation of $43,400​).

The remainder varies with sales. Both fixed and variable​ marketing, distribution, and​ customer-service costs are paid as incurred.

.

Prepare a cash budget for December

2014

and January

2015.

Supply supporting schedules for collections of​ receivables; payments for​ merchandise; and​ marketing, distribution, and​ customer-service costs.

2.

Why do Itali​'s managers prepare a cash budget in addition to the operating income​ budget?

Requirement 1. Prepare a cash budget for December2014

and January 2015.

Supply supporting schedules for collections of​ receivables; payments for​ merchandise; and​ marketing, distribution, and​ customer-service costs.

Begin by entering the appropriate amounts to calculate the collection of receviables for December2014

and January2015.

​(For amounts with a​ $0 balance, make sure to enter​ "0" in the appropriate​ cell.)

Collections of Receivables

December 2014

January 2015

From sales in:

October

$0

November

December

January

Total collections
Answer #1
1 Cash budget
December January
Beginning balance 30000 4835
Add: Collection from sales (Note:1) 503835 559470
Total cash available a 533835 564305
Less:
Cash paid for purchase (Note:2) 429940 465400
Marketing,distribution and customer service cost:
Variable costs (Note:3) 89760 97920
Fixed costs (Note:3) 9300 9300
Total payments b 529000 572620
Ending balance a-b 4835 -8315
Note:1
Collection from sales
December January
October sales
15% on the second month after purchase (287000*15%) 43050
November sales
30% by the end of month after purchase (629000*30%) 188700
15% on the second month after purchase (629000*15%) 94350
December sales
50% within 10 days (561000*50%*0.97) 272085
30% by the end of month after purchase (561000*30%) 168300
January sales
50% within 10 days (612000*50%*0.97) 296820
Total 503835 559470
Note:2
Cash paid for purchase
December January February
Sales in $ a 561000 612000 510000
Average selling price b 170 170 170
Sales in units c=a/b 3300 3600 3000
Add: Desired ending inventory 1290 1170
(570+20% of next month sales) 570+(3600*20%) 570+(3000*20%)
Total units required 4590 4770
Less: Beginning inventory 860 1290
(Inventory in $/Average unit purchase cost) (111800/130)
Units to be purchased d 3730 3480
Average unit purchase cost e 130 130
Cost of purchase f=d*e 484900 452400
Payment:
Accounts payable 139000
60% in the month 290940 271440
(484900*60%) (452400*60%)
40% in the following month 193960
(484900*40%)
Total 429940 465400
Note:3
Fixed costs =Fixed costs(incl.dep)-Depreciation=155000-43400=$ 111600
Fixed costs per month=111600/12=$ 9300
Variable costs=Total costs-Fixed costs (incl.dep)=670000-155000=$ 515000
Variable cost ratio to sales=Total variable cost for the year/Total annual sales=515000/3218750=0.16
December January
Sales a 561000 612000
Variable cost ratio to sales b 0.16 0.16
Variable costs a*b 89760 97920
2 Cash budget is prepared to ensure that company has sufficient cash to pay suppliers,employees and to pay operating expenses.
It will also enable to borrow money to meet any shortfall in cash balance.
Since,operating income budget is based on accrual concepi.It will not provide the infromation on cash level
I appreciate your ratings